[litho] Morning Litho Status 4/25
Gary J Sosa
gsosa at stanford.edu
Thu Apr 25 08:17:45 PDT 2013
Good morning All..
SVG Dev1- 2 users report wafers coming out wet
General Information/Lab Activities
Thursday, April 25, 2013
Intermittent BSA focus issue
Making arrangements for annual PM . PO Approved. Scheduling of PM’s for both aligners in progress
Does not work as standard aligner. Suspect software or configuration
Further investigation shows that system does not work in imprint mode either. EVG opened an internet session and attempted to update the machine tool configuration list. Info being passed on to HQ for further support
Bond-Align not working
PO for service approved on 4/10. Working with EVG on comprehensive plan to check out/repair his tool on the imprint tool.
Cannot change temperature
Arranging with supplier for new temperature controller
Continuous testing and investigating
1.Chillplate temp control
2. Wafer Handling
1. Replaced water pump. Testing in progress.
2. Testing various paths and wafer sizes.
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the litho